What is Molar solubility of PbBr2 (Ks= 4.67*10^-6) in 0.050M MgBr2?

The answer that was given was x= 4.59*10^-4 M

How did they find that answer?
Answer #1

MgBr2 here is Strong electrolyte

It will dissociate completely to give [Br-] = 2*[MgBr2]

= 2*0.05 M

= 0.1 M

At equilibrium:

PbBr2 <----> Pb2+ + 2 Br-

   x 0.1 + 2x

Ksp = [Pb2+][Br-]^2

4.67*10^-6=(x)*(0.1+ 2x)^2

Since Ksp is small, x can be ignored as compared to 0.1

Above expression thus becomes:

4.67*10^-6=(x)*(0.1)^2

4.67*10^-6= (x) * 10^-2

x = 4.67*10^-4 M

Answer: 4.67*10^-4 M